# Fast, Stable and Efficient Approximation of Multi-parameter Persistence Modules with MMA

### v3

@@ 4.3 Exact reconstruction with MMA
In this section, we identify the interval decomposable multi-parameter persistence modules that can be recovered exactly with MMA. Given a size parameter $\delta>0$, they correspond to modules that can decomposed into interval summands that form a subclass of the family of discretely presented interval modules, that we call the $\delta$-discretely presented interval modules.
Definition 4.4 ($\delta$-discretely presented interval module).
@@
Let $K\subseteq\mathbb{R}^{n}$ be a hyperrectangle of $\mathbb{R}^{n}$, and let $I$ be a discretely presented interval module. Given $\delta>0$, we say that $I$ is $\delta$-discretely presented in $K$ if:
@@
    (Large facets) for each point $x\in L[I]$ (resp. $U[I]$) there exists, for each facet $F$ containing $x$, an $(n-1)$-hypercube $Q^{x}_{F}$ of side length $2\delta$ such that $x\in Q^{x}_{F}$ and $Q^{x}_{F}\subseteq F$;
@@
    (Large holes) if there exists a diagonal line $l$ such that $l\cap{\rm supp}(I)=\varnothing$, then there exists an $n$-hypercube $R$ of side length $\delta$ containing ${\bf 0}$ such that for any line $l^{\prime}$ in $l+R$, one has $l^{\prime}\cap{\rm supp}(I)=\varnothing$;
@@
Assumptions 1 and 2 ensures that the pieces of ${\rm supp}(I)$ are large enough w.r.t. $\delta$, while Assumptions 3 and 4 ensure that surrounding sets of lines can detect at most one facet associated to a given codirection at a time, and that critical points of $I$ are all included in the hyperrectangle $K$ respectively.
Remark 4.5.
One might wonder whether Assumption 2 and Assumption 3 are redundant with Assumption 1. In other words, one might wonder whether it is actually possible to define an interval module with large facets and small holes, or with large facets that can share the same codirection and lie close to each other at the same time. Even though this seems to be impossible when $n=2$ (indicating that Assumption 2 and Assumption 3 might indeed be redundant with Assumption 1), it can happen when $n\geq 3$, as Figure 9 shows.
Figure 9: Example of interval module in dimension $n=3$ with large facets, small holes and some facets with the same codirection close to each other. The support of the module can be constructed by taking the (closed) red and (open) green $L$-shaped sets on (Left), and glue them together as shown in (Middle). While arbitrarily large facets can be created using this construction, the resulting interval always contains a small hole and large facets of same codirection that are close to each other. Because of this, it is possible to find a (blue) diagonal line that goes through the support without intersecting it, while lines in its surrounding set will detect some facets. (Right) View of the interval from the top showing the hole and the spatially close facets (showed in bold font). This is an example where Assumptions 1 and 4 of Definition 4.4 are satisfied, while Assumptions 2 and 3 are not.
@@
Lemma 4.6.
Let $\delta>0$ and $K$ be a hyperrectangle of $\mathbb{R}^{n}$. Let $I$ be a $\delta$-discretely presented interval module in $K$, and let $L$ be the $\delta$-grid of lines associated to $K^{2\delta}$. Then, there is a bijection between the facets of $I$ and the labels identified by Algorithm 3.
@@
Let $F$ be a facet of $L[I]$. Let $l_{F}\in L$ be a diagonal line intersecting $F$, and $b_{F}\in\mathbb{R}^{n}$ be the associated birthpoint. By Definition 4.4, item (1), there exists an $(n-1)$-hypercube $Q_{F}^{b_{F}}\subseteq F$ of side length $2\delta$ such that $b_{F}\in Q_{F}^{b_{F}}$. This ensures that for any dimension $i$ that is not in the codirection: $i\in\llbracket 1,n\rrbracket\backslash{\rm codir}(F)$, one has either $b_{F}+\delta e_{i}\in Q_{F}^{b_{F}}$ or $b_{F}-\delta e_{i}\in Q_{F}^{b_{F}}$. Since $L$ is the $\delta$-grid of lines associated to $K^{2\delta}$, and since $Q_{F}^{b_{F}}$ is an $(n-1)$-hypercube, there exists a line $l_{0}\in L$ such that $l_{F}$ belongs to the surrounding set $L_{l_{0}}$, and such that the birthpoints corresponding to the lines in $L_{l_{0}}$ are all in $Q_{F}^{b_{F}}$. This means that ${\rm codir}(F)$ is detected as a label of $b_{F}$ by Algorithm 3. Reciprocally, assume there exists a line $l_{0}\in L$ such that all birthpoints associated to the lines in the surrounding set $L_{l_{0}}$ share a coordinate along dimension $i\in\llbracket 1,n\rrbracket$, so that $i$ is a label detected by Algorithm 3. Then, the set of birthpoints $B_{L_{l_{0}}}$ has a minimal element, and thus its convex hull ${\rm conv}(B_{L_{l_{0}}})$ is in $L[I]$. Since ${\rm conv}(B_{L_{l_{0}}})$ is an $(n-1)$-hypercube of codirection $i$, it must be associated to a facet of $L[I]$ of codirection $i$ as well. The proof extends straightforwardly for deathpoints. ∎
Now that we have proved that all facets can be detected with $\delta$-grids of lines and $\delta$-discretely presented modules, we can state our first main result, which claims that it is possible to exactly recover the underlying module under the same assumptions.
Proposition 4.7 (Exact recovery).
Let $\delta>0$ and $K=R_{\alpha,\beta}$ be a hyperrectangle of $\mathbb{R}^{n}$, where $\alpha\leq_{n}\beta$. Let $I$ be a $\delta$-discretely presented interval module in $K$, and let $L$ be the $\delta$-grid of lines associated to $K^{2\delta}$. Let $C_{B}^{L}(I)$ and $C_{D}^{L}(I)$ be the $L$-birth and death corners of $I$ computed by Algorithm 4, and let $\tilde{I}=\mathrm{\mathbf{Int}}\left(\bigcup_{c\in C_{B}^{L}(I)}\bigcup_{c^{\prime}\in C_{D}^{L}(I)}R_{c,c^{\prime}}\right)$ be the approximation computed by Algorithm 2. Then, one has
${d_{\rm I}}(I,\tilde{I})={d_{\rm B}}(I,\tilde{I})=0.$ (10)
@@
As interval modules are characterized by their support, it is enough to show that $\overline{{\rm supp}(I)}=\overline{{\rm supp}(\tilde{I})}$. In the following, we thus assume that ${\rm supp}(I)$ is closed in ¯$\overline{\mathbb{R}}^{n}$. We will also use an additional definition. Let $b$ be an infinite corner computed by Algorithm 4. We say that $b^{\prime}$ is a pseudo birth corner for $b$ if:
@@
    there exists a line $l_{0}\in L$ such that
@@
Note that codirections can be extended to any finite corner (i.e., that is potentially not the pseudo corner of an infinite corner) straightforwardly, and that pseudo death corners can be defined by symmetry. We now prove Proposition 4.7.
We first show the inclusion ${\rm supp}(\tilde{I})\subseteq{\rm supp}(I)$. More specifically, we have to prove that the corners computed by Algorithm 4 all belong to ${\rm supp}(I)$. A key argument that we will use several times comes from the following lemma, which allows for a local control of the boundary of ${\rm supp}(I)$ using the hyperplanes associated to specific corners.
Lemma 4.8.
@@
Proof of (1). Let $F$ be a facet containing $b$. According to Definition 4.4, item (1), there exists an $(n-1)$-hypercube $Q^{b}_{F}$ of side length $2\delta$ such that $Q^{b}_{F}\subseteq F$ and $b\in Q^{b}_{F}$. Since $L$ is a grid, there exists a line $l\in L$ with $d_{\infty}(b,l)<\delta$ intersecting $Q^{b}_{F}$. Now, since $b\in{\rm conv}(L_{l_{0}})$, one has $d_{\infty}(l\cap H_{F},L_{l_{0}}\cap H_{F})<\delta$, where $H_{F}$ is the hyperplane containing $F$; thus, $l\in L_{l_{0}}$ (the argument is the same than in the proof of Proposition 3.13, first paragraph).
Proof of (2). By Proposition 3.13, item (2), the birthpoints associated to lines of $L_{l_{0}}$ are all contained in a ball of radius $\delta$. Thus, the unicity of the facets with given codirection comes straightforwardly from Definition 4.4, item (3).
Proof of (3). Note that the birthpoint $b$ is obviously included in the facets of $L[I]$ that contain it, which is a subset of the facets associated to the birthpoints of the lines in $L_{l_{0}}$. Now, as Lemma 4.6 ensures that the birthpoints associated to lines in $L_{l_{0}}$ are correctly labelled, the corner generated by $L_{l_{0}}$ must be on the intersection of the facets containing $b$. This ensures that
@@
Now that we have Lemma 4.8, we can prove that finite and infinite corners belong to ${\rm supp}(I)$. We will prove the results for birth corners, but the arguments for death corners are symmetric.
Finite corners. Let $b$ be a finite birth corner, associated to a set of consecutive lines $L_{l_{0}}$ for some line $l_{0}\in L$. By assumption, each birthpoint $b_{l}^{I}$, for $l\in L_{l_{0}}$, is nontrivial; and thus any birthpoint in ${\rm conv}(L_{l_{0}})$ is nontrivial as well, using Definition 4.4, item (2). Let $l\in{\rm conv}(L_{l_{0}})$ be the diagonal line passing through $b$.
Using Lemma 4.8, one has:
@@
Thus $b=b_{l}^{I}$ and $b\in{\rm supp}(I)$.
Infinite corners. Let $b$ be an infinite birth corner, and let $b^{\prime}$ be the minimal (w.r.t. $\leq_{n}$) pseudo birth corner for $b$, which is well defined by construction of $b$ (see Algorithm 4). Let $L_{l_{0}}$ be the associated set of consecutive lines $L_{l_{0}}$, for some line $l_{0}\in L$. We will show that, if $j$ is a free coordinate of $b^{\prime}$, i.e., if $j\in{\rm dir}(b^{\prime})$, then $b^{\prime}_{j}<\alpha_{j}$ (recall that $K$ is the rectangle $R_{\alpha,\beta}$). The reason we want to prove such inequalities is that they directly lead to the result. Indeed, if $b^{\prime}_{j}<\alpha_{j}$ for any $j\in{\rm dir}(b^{\prime})$, then $b^{\prime}-t\sum_{j\in{\rm dir}(b^{\prime})}e_{j}$ belongs to $L[I]$ for any $t>0$, since otherwise the line $\{b^{\prime}-t\sum_{j\in{\rm dir}(b^{\prime})}e_{j}:t>0\}$ would have to intersect a facet $F\subseteq L[I]$ of codirection $j$ for some $j\in{\rm dir}(b^{\prime})$, which would not intersect $K$, contradicting Definition 4.4, item (4). Let $j\in{\rm dir}(b^{\prime})$ be a free coordinate. By contradiction, assume that $b^{\prime}_{j}\geq\alpha_{j}$, and let $b^{j}$ denote the pseudo birth corner generated by $L_{l_{0}-\delta e_{j}}$. In particular, this means that, for any $l\in L_{l_{0}}$, $l-\delta e_{j}\in L$ and $L_{l-\delta e_{j}}\subseteq L$ since $L$ fills $K^{\delta}$. Now, if for every line $l\in L_{l_{0}}$ such that $l=l_{0}+\overrightarrow{v}$ with $\overrightarrow{v}_{j}=0$, one has that $b_{l}^{I}$ and $b_{l-\delta e_{j}}^{I}$ are on the same facets, then one has $b_{l-\delta e_{j}}^{I}=b_{l}^{I}-\delta e_{j}$, and the pseudo corner $b^{j}$ is equal to $b^{\prime}-\delta e_{j}$ by construction, as per Algorithm 4. Moreover, one has $b^{j}=b^{\prime}-\delta e_{j}\leq_{n}b^{\prime}$, contradicting the fact that $b^{\prime}$ is minimal. Hence, there is at least one line $l\in L_{l_{0}}$, $l=l_{0}+\overrightarrow{v}$ with $\overrightarrow{v}_{j}=0$, such that $b_{l}^{I}$ and $b_{l-\delta e_{j}}^{I}$ are not on the same facets, in other words, there exists a facet $F_{j}$ of $L[I]$ of codirection $j$ that intersects the (half-open) segment $[b_{l}^{I}-\delta e_{j},b_{l}^{I})$. In order to locate that facet more precisely, we will prove the following lemma:
Lemma 4.9.
@@
Without loss of generality, assume $s=0$. Since $b^{I}_{l}-te_{i}\in l-te_{i}$, it follows that $b^{I}_{l}-te_{i}$ and $b^{I}_{l-te_{i}}$ are comparable. Moreover, one must have $b^{I}_{l}-te_{i}\leq_{n}b^{I}_{l-te_{i}}$, otherwise one would have $b^{I}_{l}>_{n}b^{I}_{l}-te_{i}>_{n}b^{I}_{l-te_{i}}$, contradicting Lemma 3.9. If the points are equal, i.e., $b^{I}_{l}-te_{i}=b^{I}_{l-te_{i}}$, then one has $(b^{I}_{l})_{i}\geq(b^{I}_{l-te_{i}})_{i}$. Otherwise, if $b^{I}_{l}-te_{i}<_{n}b^{I}_{l-te_{i}}$, then
@@
Moreover, since $b^{I}_{l}$ and $b^{I}_{l-te_{i}}$ cannot be comparable as per Lemma 3.9 one must have $(b^{I}_{l-te_{i}})_{i}\leq(b^{I}_{l})_{i}$. ∎
Let $H_{j}=\left\{x\in\mathbb{R}^{n}:x_{j}=c_{j}\right\}$ be the hyperplane associated to $F_{j}$. Then, by Lemma 4.9, one has
@@
Since the lines $l$ and $l-\delta e_{j}$ both belong to the surrounding set $L_{l_{0}-\delta e_{j}}$, it follows from Lemmas 4.6, and 4.8, item (3), that ${\rm codir}(b^{j})\supseteq{\rm codir}(b^{\prime})\cup\left\{j\right\}$. Moreover, since the facets of $L[I]$ associated to ${\rm codir}(b^{j})$ are unique in a $\delta$-ball around $b^{j}$, as per Definition 4.4, item (3), they all have a unique associated value $c_{i}$ (corresponding to their associated hyperplanes). Finally, we will show that $b^{j}\leq_{n}b^{\prime}$. Let $i\in\llbracket 1,n\rrbracket$ be an arbitrary dimension.
@@
We now show that ${\rm supp}(I)\subseteq{\rm supp}(\tilde{I})$. Let $x\in{\rm supp}(I)$. We will show that there exists a birth corner $c$ such that $c\leq_{n}x$. Let $\mathcal{H}$ be the family of hyperplanes associated to the facets of $L[I]$. The corner $c$ will be defined as the limit of a sequence of points $\{x^{(k)}\}_{k\in\mathbb{N}^{*}}$in ¯$\overline{\mathbb{R}}^{n}$, defined by induction with:
@@
    • either $x^{(2)}$ is at infinity in $H_{1}$, i.e., $x^{(2)}_{j}=-\infty$ if $j\in\mathcal{J}^{1}$ and $x^{(2)}_{j}=x^{(1)}_{j}$ otherwise, and we let $c:=x^{(2)}$.
@@
    • either $x^{(k+1)}$ is at infinity in $H_{k}$, i.e., $x^{(k+1)}_{j}=-\infty$ if $j\in\mathcal{J}^{k}$ and $x^{(k+1)}_{j}=x^{(k)}_{j}$ otherwise, and we let $c:=x^{(k+1)}$.
@@
If this sequence stops at step one, i.e., $c=x^{(1)}=-\boldsymbol{\infty}$, then every birthpoint of $I$ is at $-\boldsymbol{\infty}$, the only birth corner is $c=-\boldsymbol{\infty}$, and one trivially has $c\leq_{n}x$. Hence, we assume in the following that $c$ is obtained after at least one iteration of the sequence. Note that this sequence of points has length at most $n$. Let $c^{-}$ and $c$ be the penultimate and last elements of the sequence respectively, and let $\mathcal{J}^{-}$ be the set of free coordinates associated to $c^{-}$. By construction, one has:
@@
We now show that $c$ is indeed a birth corner. If $c$ is finite, then it must belong to the intersection of $n$ hyperplanes, and it is thus a finite birth corner. Hence, we assume now that $c$ is not finite. We will construct a minimal pseudo birth corner from $c^{-}$, and show that $c$ is its associated infinite birth corner. We will consider two different cases, depending on whether $c^{-}$ is close to $K=R_{\alpha,\beta}$ or not. If $c^{-}\in K^{\delta}$, the filling property of $L$ and the size of the facets of $L[I]$ ensure that $c^{-}$ is itself a minimal pseudo birth corner, associated to $c$, which is thus an infinite birth corner. If $c^{-}\notin K^{\delta}$, then let $\overrightarrow{v}\in\mathbb{R}^{n}$ be a vector that pushes back $c^{-}$ into $K^{\delta}$, i.e., such that, for any dimension $i\in\mathcal{J}^{-}$, one has
@@
and $\overrightarrow{v}_{i}=0$ if $i\notin\mathcal{J}^{-}$. Let $S$ be the segment $[c^{-},c^{-}+\overrightarrow{v}]$. We have the two following cases:

### v4

@@ 5.2 Exact reconstruction
In this section, we identify the interval decomposable multi-parameter persistence modules that can be recovered exactly with our MMA algorithm. Given a precision parameter $\delta>0$, they correspond to modules that can decomposed into interval summands that form a subclass of the family of discretely presented interval modules, that we call the $\delta$-discretely presented interval modules.
Definition 24 ($\delta$-discretely presented interval module).
@@
Let $K\subseteq\mathbb{R}^{n}$ be a rectangle in $\mathbb{R}^{n}$, and let $k^{I}$} be a discretely presented interval module. Given $\delta>0$, we say that $I$ is $\delta$-discretely presented in $K$ if:
@@
    (Large facets) for each point $x\in L[I]$ (resp. $U[I]$) there exists, for each facet $F$ containing $x$, an (n1)(n-1)-hypercube $Q^{x}_{F}$ of side length $2\delta$ such that $x\in Q^{x}_{F}$ and $Q^{x}_{F}\subseteq F$;
@@
    (Large holes) if there exists a diagonal line $l$ such that $l\cap I=\varnothing$, then there exists an $n$-hypercube $R$ of side length $\delta$ containing $\bf 0$} such that for any line $l^{\prime}$ in $l+R$, one has $l^{\prime}\cap I=\varnothing$;
@@
Assumptions 1 and 2 ensures that the parts of $I$ are large enough w.r.t. $\delta$, while Assumptions 3 and 4 ensure that surrounding sets of lines can detect at most one facet associated to a given codirection at a time, and that critical points of $I$ are all included in the rectangle $K$, respectively.
Remark 10.
One might wonder whether Assumption 2 and Assumption 3 are redundant with Assumption 1. In other words, one might wonder whether it is actually possible to define an interval module with large facets and small holes, or with large facets that can share the same codirection and lie close to each other at the same time. Even though this seems to be impossible when $n=2$ (indicating that Assumption 2 and Assumption 3 might indeed be redundant with Assumption 1 in that case), it can happen when $n\geq 3$, as Figure 11 shows.
Figure 11: Example of interval module in dimension $n=3$ with large facets, small holes and some facets with the same codirection close to each other. The support of the module can be constructed by taking the (closed) red and (open) green $L$-shaped sets on (Left), and glue them together as shown in (Middle). While arbitrarily large facets can be created using this construction, the resulting interval always contains a small hole and large facets of same codirection that are close to each other. Because of this, it is possible to find a (blue) diagonal line that goes through the support without intersecting it, while lines in its surrounding set will detect some facets. (Right) View of the interval from the top showing the hole and the spatially close facets (showed in bold font). This is an example where Assumptions 1 and 4 of Definition 24 are satisfied, while Assumptions 2 and 3 are not.
@@
Lemma 8 (Labels are exact).
Let $\delta>0$ and $K$ be a rectangle in $\mathbb{R}^{n}$. Let $k^{I}$} be a $\delta$-discretely presented interval module in $K$, and let $L:=L_{\delta}(K^{2\delta})$ be the $\delta$-grid of lines of the offset $K^{2\delta}$. Then, there exists a bijection between the facets of $I$ and the labels identified by Algorithm 3.
@@
Let $F$ be a facet of $L[I]$. Let $l_{F}\in L$ be a diagonal line intersecting $F$, and $b_{F}\in\mathbb{R}^{n}$ be the associated birthpoint. By Definition 24, item (1), there exists an (n1)(n-1)-hypercube $Q_{F}^{b_{F}}\subseteq F$ of side length $2\delta$ such that $b_{F}\in Q_{F}^{b_{F}}$. This ensures that for any dimension $i$ that is not in the codirection: $i\in\llbracket 1,n\rrbracket\backslash{\rm codir}(F)$, one has either $b_{F}+\delta e_{i}\in Q_{F}^{b_{F}}$ or $b_{F}-\delta e_{i}\in Q_{F}^{b_{F}}$. Since $L$ is the $\delta$-grid of lines associated to $K^{2\delta}$, and since $Q_{F}^{b_{F}}$ is an (n1)(n-1)-hypercube, there exists a line $l_{0}\in L$ such that $l_{F}$ belongs to the surrounding set $L_{l_{0}}$, and such that the birthpoints corresponding to the lines in $L_{l_{0}}$ are all in $Q_{F}^{b_{F}}$. This means that ${\rm codir}(F)$ is detected as a label of $b_{F}$ by Algorithm 3.
Reciprocally, assume there exists a line $l_{0}\in L$ such that all birthpoints associated to the lines in the surrounding set $L_{l_{0}}$ share a coordinate along dimension $i\in\llbracket 1,n\rrbracket$, so that $i$ is a label detected by Algorithm 3. Then, the set of birthpoints $B_{L_{l_{0}}}$ has a minimal element, and thus its convex hull ${\rm conv}(B_{L_{l_{0}}})$ is in $L[I]$. Since ${\rm conv}(B_{L_{l_{0}}})$ is an (n1)(n-1)-hypercube of codirection $i$, it must be associated to a facet of $L[I]$ of codirection $i$ as well.
The proof extends straightforwardly for deathpoints.
Now that we have proved that all facets can be detected with $\delta$-grids of lines and $\delta$-discretely presented modules, we can state our following result, which claims that it is possible to exactly recover the underlying module under the same assumptions.
Lemma 9 (Exact recovery of intervals).
Let $\delta>0$ and $K=R_{\alpha,\beta}$ be a rectangle in $\mathbb{R}^{n}$, where $\alpha\leq_{n}\beta$. Let $k^{I}$} be a $\delta$-discretely presented interval module in $K$, and let $L:=L_{\delta}(K^{2\delta})$ be the $\delta$-grid of lines of the offset $K^{2\delta}$. Let $B=L\cap I$, and $C_{B}^{L}(B)$ and $C_{D}^{L}(B)$ be the $L$-birth and death corners of II computed by Algorithm 4, and let $\tilde{I}=\bigcup_{c\in C_{B}^{L}(B)}\bigcup_{c^{\prime}\in C_{D}^{L}(B)}R_{c,c^{\prime}}$ be the approximation computed by Algorithm 2. Then, one has:
${d_{\rm I}}\left({k^{I}},{k^{\tilde{I}}}\right)={d_{\rm B}}\left({k^{I}},{k^{\tilde{I}}}\right)=0.$ (18)
@@
As interval modules are characterized by their support, it is enough to show that $\overline{I}=\overline{\tilde{I}}$. In the following, we thus assume that $I$ is closed in $\overline{\mathbb{R}}^{n}$. We will also use an additional definition. Let $b$ be an infinite corner computed by Algorithm 4. We say that $b^{\prime}$ is a pseudo birth corner for $b$ if:
@@
    there exists a line $l_{0}\in L$ such that:
@@
Note that codirections can be extended to any finite corner (i.e., that is potentially not the pseudo corner of an infinite corner) straightforwardly, and that pseudo death corners can be defined by symmetry. We now prove Proposition 12.
We first show the inclusion ${\tilde{I}}\subseteq{I}$. More specifically, we have to prove that the corners computed by Algorithm 4 all belong to I{I}. A key argument that we will use several times comes from the following lemma, which allows for a local control of the boundary of I{I} using the hyperplanes associated to specific corners.
Lemma 10.
@@
Proof of (1). Let $F$ be a facet containing $b$. According to Definition 24, item (1), there exists an (n1)(n-1)-hypercube $Q^{b}_{F}$ of side length $2\delta$ such that $Q^{b}_{F}\subseteq F$ and $b\in Q^{b}_{F}$. Since $L$ is a grid, there exists a line $l\in L$ with $d_{\infty}(b,l)<\delta$ intersecting $Q^{b}_{F}$. Now, since $b\in{\rm conv}(L_{l_{0}})$, one has $d_{\infty}(l\cap H_{F},L_{l_{0}}\cap H_{F})<\delta$, where $H_{F}$ is the hyperplane containing $F$; thus, $l\in L_{l_{0}}$ (the argument is the same than in the proof of Proposition 7, first paragraph).
Proof of (2). By Proposition 7, item (2), the birthpoints associated to lines of $L_{l_{0}}$ are all contained in a ball of radius $\delta$. Thus, the unicity of the facets with given codirection comes straightforwardly from Definition 24, item (3).
Proof of (3). Note that the birthpoint $b$ is obviously included in the facets of $L[I]$ that contain it, which is a subset of the facets associated to the birthpoints of the lines in $L_{l_{0}}$. Now, as Lemma 8 ensures that the birthpoints associated to lines in $L_{l_{0}}$ are correctly labelled, the corner generated by $L_{l_{0}}$ must be on the intersection of the facets containing $b$. This ensures that:
@@
Now that we have Lemma 10, we can prove that finite and infinite corners belong to ${\rm supp}(I)$. We will prove the results for birth corners, but the arguments for death corners are symmetric.
Finite corners. Let $b$ be a finite birth corner, associated to a set of consecutive lines $L_{l_{0}}$ for some line $l_{0}\in L$. By assumption, each birthpoint $b_{l}^{I}$, for $l\in L_{l_{0}}$, is nontrivial; and thus any birthpoint in ${\rm conv}(L_{l_{0}})$ is nontrivial as well, using Definition 24, item (2). Let $l\in{\rm conv}(L_{l_{0}})$ be the diagonal line passing through $b$.
Using Lemma 10, one has:
@@
Thus $b=b_{l}^{I}$ and $b\in{I}$.
Infinite corners. Let $b$ be an infinite birth corner, and let $b^{\prime}$ be the minimal (w.r.t. $\leq_{n}$) pseudo birth corner for $b$, which is well defined by construction of $b$ (see Algorithm 4). Let $L_{l_{0}}$ be the associated set of consecutive lines $L_{l_{0}}$, for some line $l_{0}\in L$. We will show that, if $j$ is a free coordinate of $b^{\prime}$, i.e., if $j\in{\rm dir}(b^{\prime})$, then $b^{\prime}_{j}<\alpha_{j}$ (recall that $K$ is the rectangle $R_{\alpha,\beta}$). The reason we want to prove such inequalities is that they directly lead to the result. Indeed, if $b^{\prime}_{j}<\alpha_{j}$ for any $j\in{\rm dir}(b^{\prime})$, then $b^{\prime}-t\sum_{j\in{\rm dir}(b^{\prime})}e_{j}$ belongs to $L[I]$ for any $t>0$, since otherwise the line $\{b^{\prime}-t\sum_{j\in{\rm dir}(b^{\prime})}e_{j}:t>0\}$ would have to intersect a facet $F\subseteq L[I]$ of codirection $j$ for some $j\in{\rm dir}(b^{\prime})$, which would not intersect $K$, contradicting Definition 24, item (4).
Let $j\in{\rm dir}(b^{\prime})$ be a free coordinate. By contradiction, assume that $b^{\prime}_{j}\geq\alpha_{j}$, and let $b^{j}$ denote the pseudo birth corner generated by $L_{l_{0}-\delta e_{j}}$. In particular, this means that, for any $l\in L_{l_{0}}$, $l-\delta e_{j}\in L$ and $L_{l-\delta e_{j}}\subseteq L$ since $L$ fills $K^{\delta}$. Now, if for every line $l\in L_{l_{0}}$ such that $l=l_{0}+\overrightarrow{v}$ with $\overrightarrow{v}_{j}=0$, one has that $b_{l}^{I}$ and $b_{l-\delta e_{j}}^{I}$ are on the same facets, then one has $b_{l-\delta e_{j}}^{I}=b_{l}^{I}-\delta e_{j}$, and the pseudo corner $b^{j}$ is equal to $b^{\prime}-\delta e_{j}$ by construction, as per Algorithm 4. Moreover, one has $b^{j}=b^{\prime}-\delta e_{j}\leq_{n}b^{\prime}$, contradicting the fact that $b^{\prime}$ is minimal. Hence, there is at least one line $\in L_{l_{0}}, l=l$$l_{0}+\overrightarrow{v}$ with $\overrightarrow{v}_{j}=0$, such that $b_{l}^{I}$ and $b_{l-\delta e_{j}}^{I}$ are not on the same facets, in other words, there exists a facet $F_{j}$ of $L[I]$ of codirection $j$ that intersects the (half-open) segment $b_{l}^{I}-\delta e_{j},b_{l}^{I})$. In order to locate that facet more precisely, we will prove the following lemma.
Lemma 11.
@@
Without loss of generality, assume $s=0$. Since $b^{I}_{l}-te_{i}\in l-te_{i}$, it follows that $b^{I}_{l}-te_{i}$ and $b^{I}_{l-te_{i}}$ are comparable. Moreover, one must have $b^{I}_{l}-te_{i}\leq_{n}b^{I}_{l-te_{i}}$, otherwise one would have $b^{I}_{l}>_{n}b^{I}_{l}-te_{i}>_{n}b^{I}_{l-te_{i}}$, contradicting Lemma 1. If the points are equal, i.e., $b^{I}_{l}-te_{i}=b^{I}_{l-te_{i}}$, then one has $(b^{I}_{l})_{i}\geq(b^{I}_{l-te_{i}})_{i}$. Otherwise, if $b^{I}_{l}-te_{i}<_{n}b^{I}_{l-te_{i}}$, then:
@@
Moreover, since $b^{I}_{l}$ and $b^{I}_{l-te_{i}}$ cannot be comparable as per Lemma 1 one must have $(b^{I}_{l-te_{i}})_{i}\leq(b^{I}_{l})_{i}$.
Let $H_{j}=\left\{x\in\mathbb{R}^{n}:x_{j}=c_{j}\right\}$ be the hyperplane associated to $F_{j}$. Then, by Lemma 11, one has:
@@
Since the lines $l$ and $l-\delta e_{j}$ both belong to the surrounding set $L_{l_{0}-\delta e_{j}}$, it follows from Lemmas 8, and 10, item (3), that ${\rm codir}(b^{j})\supseteq{\rm codir}(b^{\prime})\cup\left\{j\right\}$. Moreover, since the facets of $L[I]$ associated to ${\rm codir}(b^{j})$ are unique in a $\delta$-ball around $b^{j}$, as per Definition 24, item (3), they all have a unique associated value $c_{i}$ (corresponding to their associated hyperplanes).
Finally, we will show that $b^{j}\leq_{n}b^{\prime}$. Let $i\in\llbracket 1,n\rrbracket$ be an arbitrary dimension.
@@
We now show that ${I}\subseteq{\tilde{I}}$. Let $x\in{I}$. We will show that there exists a birth corner $c$ such that $c\leq_{n}x$. Let $\mathcal{H}$ be the family of hyperplanes associated to the facets of $L[I]$. The corner $c$ will be defined as the limit of a sequence of points $\{x^{(k)}\}_{k\in\mathbb{N}^{*}}$in $\overline{\mathbb{R}}^{n}$, defined by induction with:
@@
    • either $x^{(2)}$ is at infinity in $H_{1}$, i.e., $x^{(2)}_{j}$$\infty$ if $j\in\mathcal{J}^{1}$ and $x^{(2)}_{j}=x^{(1)}_{j}$ otherwise, and we let $c:=x^{(2)}$.
@@
    • either $x^{(k+1)}$ is at infinity in $H_{k}$, i.e., $x^{(k+1)}_{j}$$\infty$ if $j\in\mathcal{J}^{k}$ and $x^{(k+1)}_{j}=x^{(k)}_{j}$ otherwise, and we let $c:=x^{(k+1)}$.
@@
[…]
@@
We now show that $c$ is indeed a birth corner. If $c$ is finite, then it must belong to the intersection of $n$ hyperplanes, and it is thus a finite birth corner. Hence, we assume now that $c$ is not finite. We will construct a minimal pseudo birth corner from $c^{-}$, and show that $c$ is its associated infinite birth corner. We will consider two different cases, depending on whether $c^{-}$ is close to $K=R_{\alpha,\beta}$ or not. If $c^{-}\in K^{\delta}$, the filling property of $L$ and the size of the facets of $L[I]$ ensure that $c^{-}$ is itself a minimal pseudo birth corner, associated to $c$, which is thus an infinite birth corner. If $c^{-}\notin K^{\delta}$, then let $\overrightarrow{v}\in\mathbb{R}^{n}$ be a vector that pushes back $c^{-}$ into $K^{\delta}$, i.e., such that, for any dimension $i\in\mathcal{J}^{-}$, one has:
@@
and $\overrightarrow{v}_{i}=0$ if $i\notin\mathcal{J}^{-}$. Let $S$ be the segment $c^{-},c^{-}+\overrightarrow{v}]$. We have the two following cases: